OpenStudy (anonymous):

What is the solution to the equation 4x = 42? x = 6 x = 10 x = 10.5 x = 10.6

OpenStudy (katrinakaif):

Your goal is to bring x alone. To do so, we have to preform the opposite of multiplication, which is utterly division. Divide 4 on both sides and see what you get as x.

OpenStudy (anonymous):

\[4x=42\\ x=42/4\\ x=21/2\\ x=10.5 \]